how would you fiqure out the order pair of 5x-2y=4 when x= 4,-2,0,2
Answer by stanbon(75887)

how would you fiqure out the order pair of
5x-2y=4 when x= 4,-2,0,2
---
Solve for "y":
y = (5/2)x-2
---
Then:
(4,8)
(-2,-7)
(0,-2)
(2,3)
==========
Cheers,
Stan H.